# DreamSim: Learning New Dimensions of Human Visual Similarity using Synthetic Data

### v1

@@ B.2 Additional Visualizations
Qualitative metric comparisons. In Sec. 5.2 and Fig. 67 of the main text we quantitatively analyze the differences between metrics. Here, we also provide a qualitative analysis by comparing image pairs that achieve high and low similarity scores for each metric.
@@
In Fig. 15 we plot, for each image pair in our dataset, the similarity scores from our metric against similarity scores from DINO, OpenCLIP, DISTS, and LPIPS. We show the same for image pairs drawn from the COCO dataset. In Fig. 16 and Fig. 17 we show the pairs where our metric most agrees and most disagrees with DINO, OpenCLIP, DISTS, and LPIPS, for our dataset’s test set and the COCO dataset. Note that for COCO we draw random pairs of images that share at least instance category so that pairs have some semantic commonality, enabling better visualization of qualitative differences between metrics.
@@
Attention Map Visualizations. As an alternate way of understanding our model’s similarity decisions, we visualize the transformer attention maps following Chefer et al. chefer2021transformer . Our model consists of finetuned versions of DINO, CLIP, and OpenCLIP backbones, and the resulting attention map places the largest activations on the finetuned DINO backbone (Fig. 18-left). Accordingly, the overall attention map looks largely similar to the attention map constructed from only the finetuned DINO branch. Compared to the pretrained versions of each model backbone, the finetuned model better captures full object identity (such as over the entire body of the fish), while also minimizing spurious attention values in the background (Fig. 18-right). Consistent with earlier analysis, this supports the fact that the foreground plays a larger role in the DreamSim similarily judgement than the background.
@@
As the DINO model has the largest contribution in the attention maps, Fig. 19 shows additional examples focusing on the difference between the finetuned DINO backbone within our model, and the pretrained DINO backbone prior to finetuning. This visualizes how the DINO backbone changes as it is finetuned on our dataset. Our finetuned model better captures the foreground object, while attention maps from the pretrained DINO backbone may only focus on small portions of the object. In the lobster example (Fig. 19 top-right), our model places attention on the relevant parts of the object, such as the lobster body rather than the claws in the A image as the claws do not appear in the other two images.
B.3 Additional Results on Applications
Guided Diffusion Details. In Fig. 9 we show examples of using our metric to guide a $256\times 256$ unconditional ImageNet-trained diffusion model NEURIPS2021_49ad23d1 using an input prompt image. At every training step we take the spherical distance loss between our model’s embeddings of 64 random crops of the input image and synthesized image. We add this loss to the L2 Total Variation loss Mahendran2014UnderstandingDI of the synthesized image, following the procedures of mullisClipGuided ; kcrowsonKDiffusion .
User study on image retrieval. Please see additional results on image retrieval and image reconstruction in the attached HTML page. To quantitatively evaluate our image retrieval results, we conduct a user study to collect preferences across returned results from LPIPS, DISTS, DINO, OpenCLIP, and DreamSim. In Figure 20, we provide a detailed breakdown of user preferences for neighbors 1 through 10 across ImageNet-R and COCO, along with error bars marking 1 standard deviation above and below each result.
Figure 20: User preferences for image retrieval results by metric. We conduct a user study that collects preferences for retrieval results output by LPIPS, DISTS, DINO, OpenCLIP, and DreamSim. For most instances of the first 10 nearest neighbors, users preferred our metric’s output, followed by DINO and OpenCLIP. We visualize one standard deviation above and below each bar.
@@ Appendix C Discussion
Broader Impacts and Limitations. Our model and dataset are developed from pretrained Stable Diffusion, CLIP, OpenCLIP and DINO backbones. As such, our model can inherit and propagate biases existing in these models for decisions on downstream tasks. Our dataset is generated using prompts that describe a single-word category and is filtered to remove images containing sensitive topics, such as violence or human faces. As a result, the dataset has a large focus on object-centric domains, and content containing humans is considered out-of-domain. The resulting dataset and model does not capture the full range of human similarity judgements, but only the variations that we can capture in our synthetically-generated dataset.
Licenses. The icons for the teaser figure as well as the images for the inversion experiments are licensed by Adobe Stock, under the Adobe Stock Standard License, and by Pixabay, under their content license.

### v3

@@ B.2 Additional Visualizations
Qualitative metric comparisons. In Section 5.2 and Figure 67 of the main text we quantitatively analyze the differences between metrics. Here, we also provide a qualitative analysis by comparing image pairs that achieve high and low similarity scores for each metric.
@@
In Figure 19 we plot, for each image pair in our dataset, the similarity scores from our metric against similarity scores from DINO, OpenCLIP, DISTS, and LPIPS. We show the same for image pairs drawn from the COCO dataset. In Figure 20 and Figure 21 we show the pairs where our metric most agrees and most disagrees with DINO, OpenCLIP, DISTS, and LPIPS, for our test set and the COCO dataset. Note that for COCO we draw random pairs of images that share at least instance category so that pairs have some semantic commonality, enabling better visualization of qualitative differences between metrics.
@@
Attention Map Visualizations. As an alternate way of understanding our model’s similarity decisions, we visualize the transformer attention maps following Chefer et al. chefer2021transformer . Our model consists of finetuned versions of DINO, CLIP, and OpenCLIP backbones, and the resulting attention map places the largest activations on the finetuned DINO backbone (Figure 22 (left)). Accordingly, the overall attention map looks largely similar to the attention map constructed from only the finetuned DINO branch. Compared to the pretrained versions of each model backbone, the finetuned model better captures full object identity (such as over the entire body of the fish), while also minimizing spurious attention values in the background (Figure 22 (right)). Consistent with earlier analysis, this supports the fact that the foreground plays a larger role in the DreamSim similarily judgment than the background.
@@
As the DINO model has the largest contribution in the attention maps, Figure 23 shows additional examples focusing on the difference between the finetuned DINO backbone within our model, and the pretrained DINO backbone prior to finetuning. This visualizes how the DINO backbone changes as it is finetuned on our dataset. Our finetuned model better captures the foreground object, while attention maps from the pretrained DINO backbone may only focus on small portions of the object. In the lobster example (Figure  23 (top-right)), our model places attention on the relevant parts of the object, such as the lobster body rather than the claws in the A image as the claws do not appear in the other two images.
@@ Appendix C Discussion
Broader Impacts and Limitations. Our model and dataset are developed from pretrained Stable Diffusion, CLIP, OpenCLIP and DINO backbones. As such, our model can inherit and propagate biases existing in these models for decisions on downstream tasks. Our dataset is generated using prompts that describe a single category, and we manually filter out categories that may generate sensitive content, such as violence or malformed human faces. As a result, the dataset has a large focus on object-centric domains, and content containing humans is considered out-of-domain. The resulting dataset and model does not capture the full range of human similarity judgements, but only the variations that we can capture in our synthetically-generated dataset.
Licenses. The icons for Figure 1 as well as the images for the inversion experiments are licensed by Adobe Stock, under the Adobe Stock Standard License, and by Pixabay, under their content license.